verify identinty
sec^4x – (tan^4x + sec^2x) = sec^2x · sin^2x

Answer by Alan3354(69443)

sec^4x – (tan^4x + sec^2x) = sec^2x · sin^2x
sec^4x – tan^4x - sec^2x) = sec^2x*sin^2x = sin^2/cos^2
Multiply thru by cos^4 and change to sines & cosines
1 - sin^4 - cos^2 = sin^2cos^2
sin^2 - sin^4 = sin^2cos^2
Divide by sin^2
1 - sin^2 = cos^2
--------
Always change everything to sines and cosines.